1. Calculate the pH of the acetate buffer, mmol. HA, mmol. A-, and mmol. H3O+ after 4mL, 8mL, 12mL, 16mL, 20mL, 24mL, and 28mL of 0.100M HCl is added to the acetate buffer.

2. Calculate the pH of the acetate buffer, mmol. HA, mmol. A-, and mmol. H3O+ after 4mL, 8mL, 12mL, 16mL, 20mL, 24mL, and 28mL of 0.100M NaOH is added to the acetate buffer.

3. Calculate the pH of the ammonia buffer, mmol. HA, mmol. A-, and mmol. H3O+ after 4mL, 8mL, 12mL, 16mL, 20mL, 24mL, and 28mL of 0.100M HCl is added to the ammonia buffer.

4. Calculate the pH of the ammonia buffer, mmol. HA, mmol. A-, and mmol. H3O+ after 4mL, 8mL, 12mL, 16mL, 20mL, 24mL, and 28mL of 0.100M NaOH is added to the ammonia buffer.

Given:

0.160M NaC2H3O2 / 0.200M HC2H3O2

0.160M NH4Cl / 0.200M NH3

0.100M NaOH

0.100M HCl

Volume of Buffer: started by adding 10.00mL buffer with 30.0 mL distilled water

Please help! I know it is a lot of calculations, but even if you could just show my the steps for a few of the questions it would be a great help!
Answer #1

for acidic Buffer, pt= pka + log [salt] land] sme c we have [Chorona 3 = o.16OM P [ CHz Cook] = 0.200M 0.000 form : P = 4.74me o.1 (M) He contains 8 mt xo.lM 1000m - 0.0008 mol Ations. 0.0008 mot Ht combines with 0.0088 mer che coon to form 0.0008 mNO @ pH of the acetate buffer = 4.64. when 4 mt 0.100 M Naot added. Then ho.cf moles 4 x 0.1 1000 1 = = 0.0ooumor. 0 OH 0.0not moles of its coolt = 0.0080 - 0.0028 -0.0032 mots. soles pH = 4.74 + log 0.00012 = 4.99 0.00 52 *Cook mots pt of a amoniao pH of ammonia buffer= 9.35. Now. 4 ml oil (M) NaOH added to the buffer solution. 4 ml xoll :. No. of moles of ot-added - 10